# Triplet Interaction Improves Graph Transformers: Accurate Molecular Graph Learning with Triplet Graph Transformers

### v1

@@ B.1 Axial Attention
GEM-2 (Liu et al., 2022a) introduced axial attention, which can be simplified when only we consider pairwise channels as follows:
where is the value vector of the pair and is the attention weight between pairs and (we have re-used the notations from node-to-node attention for consistency). and are the query and key vectors of pairs and respectively. This is a generalization of the self-attention mechanism for pairs and has a computational complexity of . However, notice that the third arm of the triangle, i.e., the pair is not considered in this process. GEM-2 (Liu et al., 2022a) instead uses a third-order channel to provide positional information to this attention process. However, in the absence of third-order information like angles, axial attention doesn’t perform so well. Another update is done by changing the direction of the aggregated pairs, i.e., from to , with the weights .
@@ B.2 Triangular Update
The triangular update proposed by Alphafold (Jumper et al., 2021) and later adopted by UniMol+ (Lu et al., 2023) takes the form:
@@
Notice that, in this case, the update is a simple matrix/tensor multiplication which can have subcubic complexity. However, the information flow from pair to pair is mediated only by the third arm of the triangle, i.e., the pair . Unlike a true attention process, cannot “select” which pairs to aggregate. The information passed for each set is a scalar, which means that many sets are required compared to a few attention heads. Also, this summation/aggregation process is unbounded which can be problematic if the input graphs vary in size drastically.
@@ Appendix C The Edge-augmented Graph Transformer (EGT)
where and are the query and key vectors of nodes and respectively, is the dimension of the key vectors, is the output vector which is used to update the node embedding , and is the scaled dot product between the query and key vectors. The edge channels participate in the attention mechanism by (i) providing a bias term and (ii) providing a gating term which passes through a sigmoid function to directly control the flow of information from node to node . Both the bias and gating terms are computed from projections of the edge embeddings . This is done for each head of the multi-head attention mechanism in each layer. The node channels are updated from the projection of the concatenation of of all heads. The edge channels are also updated from a projection of the concatenation of the scaled dot product of all heads. This way, EGT ensures two-way communication between the node and edge channels in the multi-head attention layer. This is in contrast to architectures like UniMol+ (Lu et al., 2023) where the edge channels are updated from an outer product of the node embeddings which adds additional computational cost.
@@ E.1 PCQM4Mv2
Despite our models having a higher parameter count compared to the previous SOTA UniMol+, when combining parameters of the distance predictor and the gap predictor, it’s crucial to recognize that direct parameter count comparisons can be misleading, especially for iterative models like UniMol+, where parameters are shared across iterations, contrasting with non-iterative models like ours. To illustrate this we train a TGT-Ag model where the consequent layers share the same set of parameters, dubbed as TGT-Agx2 reducing parameter count by half yet still outperforming UniMol+. However, we do not resort to this form of parameter sharing because although it makes the model parameter-efficient, it does not significantly reduce the computational and memory costs of training and inference. Instead, we focus on compute efficiency, i.e., to get the best possible result for a given compute budget.
@@ E.3 QM9
To highlight the transfer of knowledge to related quantum chemical prediction tasks we take the pretrained task predictor from our PCQM4Mv2 experiment and finetune it on the QM9 (Ramakrishnan et al., 2014) dataset. QM9 is a quantum chemistry benchmark consisting of 134k small organic molecules. The ground truth 3D coordinates are provided on this dataset which can be used during inference, so the distance predictor is not required. The task instead is to predict different Quantum Mechanical properties as accurately as possible from the given 3D graph. We report finetuning performance on a subset of 7 tasks from the 12 tasks in QM9, namely, dipole moment (), isotropic polarizability (), HOMO (), LUMO () energies and their difference (), Zero Point Vibrational Energy (ZPVE) and Heat Capacity (). The results are presented in terms of MAE. Energies are expressed in meV. We use the same dataset splitting strategy as Transformer-M (Luo et al., 2022) to form validation and test splits consisting of 10,000 and 10,831 molecules respectively. We use MAE loss (normalized by the standard deviation of the task) and the Adam optimizer to finetune the pretrained task predictor model.
@@ Appendix F Hyperparameters
The hyperparameters used for each dataset are presented in Table 11. For PCQM4Mv2 and OC20 we list the hyperparameters for both the distance and the task predictor models and both training and finetuning. For QM9, we only list the hyperparameters for finetuning. For MOLPCBA, LIT-PCBA, and TSP we only show the hyperparameters for training from scratch. The missing hyperparameters do not apply to the corresponding dataset or model. For QM9 no secondary distance denoising objective is used. For LIT-PCBA, 0 triplet interaction heads indicate that an EGT is used without any triplet interaction module. For TSP datasets we train two models with 4 and 16 layers for parameter budgets 100K and 500K respectively, which otherwise use the same hyperparameters.
@@ Appendix G Additional Results
OC20 IS2RE The breakdown of performance on the OC20 IS2RE validation and test results are presented in Table 12 and Table 13 respectively over the dataset splits ID (In Domain) and OOD (Out Of Domain) Adsorbates, Catalyst and Both. Notice that, TGT-At and UniMol+ outperform all other models for all splits.

### v2

@@ B.1 Axial Attention
GEM-2 (Liu et al., 2022a) introduced axial attention, which can be simplified when we consider only pairwise channels as follows:
where is the value vector of the pair and is the attention weight between pairs and (we have re-used the notations from node-to-node attention for consistency). and are the query and key vectors of pairs and , respectively. This is a generalization of the self-attention mechanism for pairs and has a computational complexity of . However, notice that the neighboring pair of the 3-tuple is not considered in this process. GEM-2 (Liu et al., 2022a) instead uses a 3rd-order channel to provide positional information to this attention process. However, in the absence of 3rd-order information like angles, axial attention does not perform well. Another update is done by changing the direction of the aggregated pairs, i.e., from to , with the weights .
@@ B.2 Triangular Update
The triangular update proposed by AlphaFold 2 (Jumper et al., 2021) and later adopted by UniMol+ (Lu et al., 2023) takes the form:
@@
Notice that, in this case, the update is a simple matrix/tensor multiplication which can have subcubic complexity. However, the information flow from pair to pair is mediated only by the pair . Unlike a true attention process, cannot “select” which pairs to aggregate. The information passed for each set is a scalar, which means that many sets are required compared to a few attention heads. Also, this summation/aggregation process is unbounded which can be problematic if the input graphs vary in size drastically.
@@ Appendix C The Edge-augmented Graph Transformer (EGT)
where and are the query and key vectors of nodes and , respectively, is the dimension of the key vectors, and is the output vector which is used to update the node embedding . The attention logit is the summation of the scaled dot product (between the query and key vectors) and the attention bias. The edge channels participate in the attention mechanism by (i) providing a bias term and (ii) providing a gating term which passes through a sigmoid function to directly control the flow of information from node to node . Both the bias and gating terms are computed from projections of the edge embeddings . This is done for each head of the multi-head attention mechanism in each layer. The node channels are updated from the projection of the concatenation of of all heads. The edge channels are also updated from a projection of the concatenation of the attention logits of all heads. This way, EGT ensures two-way communication between the node and edge channels in the multi-head attention layer. This is in contrast to architectures like UniMol+ (Lu et al., 2023) where the edge channels are updated from an outer product of the node embeddings which adds additional computational cost.
@@ E.1 PCQM4Mv2
Despite our models having a higher parameter count compared to the previous SOTA UniMol+, when combining parameters of the distance predictor and the gap predictor, it is crucial to recognize that direct parameter count comparisons can be misleading, especially for iterative models like UniMol+, where parameters are shared across iterations, contrasting with non-iterative models like ours. To illustrate this we train a TGT-Ag model where the consequent layers share the same set of parameters, dubbed as TGT-Agx2 reducing parameter count by half yet still outperforming UniMol+. However, we do not resort to this form of parameter sharing because although it makes the model parameter-efficient, it does not significantly reduce the computational and memory costs of training and inference. Instead, we focus on compute efficiency, i.e., to get the best possible result for a given compute budget.
@@ E.3 QM9
To highlight the transfer of knowledge to related quantum chemical prediction tasks we take the pretrained task predictor from our PCQM4Mv2 experiment and finetune it on the QM9 (Ramakrishnan et al., 2014) dataset. QM9 is a quantum chemistry benchmark consisting of 134k small organic molecules. The ground truth 3D coordinates are provided on this dataset which can be used during inference, so the distance predictor is not required. The task instead is to predict different Quantum Mechanical properties as accurately as possible from the given 3D graph. We report finetuning performance on a subset of 7 tasks from the 12 tasks in QM9, namely, dipole moment (), isotropic polarizability (), HOMO (), LUMO () energies and their difference (), Zero Point Vibrational Energy (ZPVE) and Heat Capacity (). The results are presented in terms of MAE. Energies are expressed in meV. We use the same dataset splitting strategy as Transformer-M (Luo et al., 2022) to form validation and test splits consisting of 10,000 and 10,831 molecules, respectively. We use MAE loss (normalized by the standard deviation of the task) and the Adam optimizer to finetune the pretrained task predictor model.
@@ Appendix F Hyperparameters
The hyperparameters used for each dataset are presented in Table 11. For PCQM4Mv2 and OC20 we list the hyperparameters for both the distance and the task predictor models and both training and finetuning. For QM9, we only list the hyperparameters for finetuning. For MOLPCBA, LIT-PCBA, and TSP we only show the hyperparameters for training from scratch. The missing hyperparameters do not apply to the corresponding dataset or model. For QM9 no secondary distance denoising objective is used. For LIT-PCBA, 0 triplet interaction heads indicate that an EGT is used without any triplet interaction module. For TSP datasets we train two models with 4 and 16 layers for parameter budgets 100K and 500K, respectively, which otherwise use the same hyperparameters.
@@ Appendix G Additional Results
OC20 IS2RE The breakdown of performance on the OC20 IS2RE validation and test results are presented in Table 12 and Table 13, respectively over the dataset splits ID (In Domain) and OOD (Out Of Domain) Adsorbates, Catalyst and Both. Notice that, TGT-At and UniMol+ outperform all other models for all splits.